OpenStudy (anonymous):

how to find the angle? tan x + 4tan x = 4sec x

OpenStudy (anonymous):

Are any options given? Because I'm not sure of my answer.

OpenStudy (anonymous):

the answer is 41.81 and 138.19 in degree form

OpenStudy (king):

secx=5/3?

OpenStudy (king):

tanx=4/3?

OpenStudy (king):

so the triangle is a right triangle with sides 3,4,5

OpenStudy (anonymous):

how did u get 5/3 and 4/3?

OpenStudy (anonymous):

Sin = Opposite/hypotenuse cos= adjacent/hypotenuse tan = oppoite/adjacent

OpenStudy (anonymous):

With this when you simplify the expression you will get sinx=4/5, which is opposite/hypotenuse. So, by pythagoras theorem, the remaining side is 3

OpenStudy (king):

umm.see simplify the given equation we get 5tanx=4secx tanx/secx=4/5 squaring on both sides tan^2x/sec^2x=16/25 we know that acc. to trig identity tan^2x=sec^2x-1 so, (sec^2x-1)/sec^2x=16/25 25sec^2x-25=16sec^2x 9sec^2x=25 root on both sides secx=5/3 substitute this in first equation to get tanx=4/3

OpenStudy (king):

yup even abdul's method is rite!

OpenStudy (king):

and yeah its easier!but i thought that in a exam,my teacher doesn't giv me marks if i dont use identities so teahcers might cut marks so i told u this method!but abdul's way is easier!

OpenStudy (anonymous):

this is the first tyme i saw the method...thanks..

OpenStudy (anonymous):

There are many ways in doing maths, we need to get the solution, the easier the better. That shouldn't affect the marks.